An atom in an excited state 2.75 eV above the ground state remains in that excited state 2.05 µs before moving to the ground state.

(a) Find the frequency of the emitted photon.
THz

(b) Find the wavelength of the emitted photon.
nm

(c) Find the approximate uncertainty in energy of the photon.
ΔE ≥  peV
Answer #1

a) Given that E = 2.75 ev ot = 2:05ms 22.05x10-6 : Energy Esho Frequency o= E. | has Plancks constant = 6.675 X 10-345-5 0 2.A E - c) By Heisenbergs uncertainity principle AE. At = ni 41 AE = 5 - 415x at DE = 6.625 *10-34 4T X2:05xlo -6. = 6.625x10-